OpenStudy (anonymous):

If p∨q is true, then what must be true about the truth values of p and q? Where p and q are statements, "p and q" is false if p is false. Which of the following statements is the contrapositive of "If I have money, then I spend it"? If I don't have money, then I don't spend it. If I spend money, then I have it. If I don't spend money, then I don't have it.

OpenStudy (ghostgate):

If p V q is true also known as p or q it would relate to the Disjunction truth table. The values of p and q that equal true outcomes for p V q are: p = True - q = True equals True p = True - q = false equals True p = False - q = True equals True However if there both false than the outcome is false. p and q truth table also known as p ^ q is: p = True and q = True equals True but if either of these changes to false the equivalent is false. So your second statement or question is correct. Now a contrapositive is not only changing the locations of "p" and "q", but also negating them as well. For example: p = true and q = false Contrapositive is: q = true and p = false Get it? So your final question's answer would be: C - If I don't spend money, then I don't have it. Hope this helps, G'day.